OpenStudy (anonymous):

write the slopeintercept equation for the line that passes through(-7,6) and is perpindicular to -7x+9y=-2 plz help

OpenStudy (mathmate):

Can you find the slope of -7x+9y=-2 ?

OpenStudy (anonymous):

i do not understand this at all

OpenStudy (mathmate):

Do you want to know how to solve the problem, or do you just want the answer?

OpenStudy (anonymous):

i wana know how to do the problem but i need the answer im on a time crunch

OpenStudy (mathmate):

Will you continue to work on how to do it after I give you the answer?

OpenStudy (anonymous):

well we can wk it now cause i have to know how to do it,if i cant get right can u help with answer

OpenStudy (mathmate):

Sure! First, do you know how to change the equation to the slope-intercept form?

OpenStudy (anonymous):

i know nothing at all

OpenStudy (anonymous):

when i read about it i dont comprehend it

OpenStudy (mathmate):

Do you take math classes, or are you home schooled?

OpenStudy (anonymous):

im in classes in college with 4.0 but this math is killing it...worst subject :(

OpenStudy (mathmate):

Looks like you need to put more time on this math thing. The equation represents a line in a graph, is that OK?

OpenStudy (anonymous):

k

OpenStudy (mathmate):

Like this: 

OpenStudy (anonymous):

its my only class at moment an im all into it i study alot just not grasping it

OpenStudy (mathmate):

That's OK, we have to start somewhere! The equation can be put into different forms, for example: -7x+9y=-2 If we add 7x to both sides, the equation becomes 9y=7x-2 Now if we divide by 9 on both sides, it becomes y=(7/9)x - (2/9) Is that ok so far?

OpenStudy (anonymous):

yes just lookin it over

OpenStudy (mathmate):

The equation has now been transformed into the "slope-intercept" form, which has y on the left, and two terms on the right (with x and without x).

OpenStudy (anonymous):

ok so far so good

OpenStudy (mathmate):

Great! The number to which "x" is multiplied (7/9) is called the slope of the line. We usually represent this number by the letter m. The bigger this number, the steeper the line. A 45 degree line has a slope of 1, and a slope of 100 is almost vertical, and a slope of 0 is horizontal.